OpenStudy (anonymous):

Rationalize the denominator of 3/(4+sqrt5)+(3-2sqrt5)

OpenStudy (earthcitizen):

it's not quite clear, do you mean, \[3/((4+\sqrt{5})+(3-2\sqrt{5}))\]

OpenStudy (anonymous):

If you mean what EarthCitizen believes you do Then the denominator can be simplified as \[1-\sqrt{5}\] If you multiply by the complex conjugate \[(1+\sqrt{5})(1-\sqrt5)\] (1-5) -4 You also have to multiply the numerator, so \[3+3\sqrt{5}\] Divide the whole thing by -4, and get \[(3+3\sqrt{5})/-4\]

OpenStudy (anonymous):

My options are [3\sqrt{5}\-5+7sqrt{5}] \[21+3\sqrt{5}/44\] \[21+\sqrt{5}/44\] \[21-3\sqrt{5}/54\]

OpenStudy (anonymous):

Ah, that means I did something wrong. Let me check my work.

OpenStudy (anonymous):

All right. It's the second option, 21+3√5 /44 When first simplifying the denominator, 4+3 is 7, meaning that all the ones should be sevens.

OpenStudy (anonymous):

okay thank you